# Bias and Fairness in Large Language Models: A Survey

### v2

@@ 5.1.5 Projection-based Mitigation
Ravfogel et al. (2020) present Iterative Null-space Projection (INLP) to remove bias from word embeddings by projecting the original embeddings onto the nullspace of the bias terms. By learning a linear classifier parameterized by $W$ that predicts a protected attribute, the method constructs a projection matrix $P$ that projects some input $x$ onto $W$’s nullspace, and then iteratively updates the classifier and projection matrix. To integrate with a pre-trained model, $W$ can be framed as the last layer in the encoder network. Adapting INLP to a non-linear classifier, Iskander et al. (2023) proposes Iterative Gradient-Based Projection (IGBP), which leverages the gradients of a neural protected attribute classifier to project representations to the classifier’s class boundary, which should make the representations indistinguishable with respect to the protected attribute. Liang et al. (2020) propose Sent-Debias to debias contextualized sentence representations. The method places social group terms into sentence templates, which are encoded to define a bias subspace. Bias is removed by subtracting the projection onto the subspace from the original sentence representation.
However, removing the concept of gender or any other protected attribute altogether may be too aggressive and eliminate important semantic or grammatical information. To address this, Limisiewicz & Mareček (2022) distinguish a gender bias subspace from the embedding space, without diminishing the semantic information contained in gendered words like pronouns. They use an orthogonal transformation to probe for gender information, and discard latent dimensions corresponding to bias, while keeping dimensions containing grammatical gender information. In their method OSCAR, Dev et al. (2021) also perform less-aggressive bias removal to maintain relevant semantic information. They orthogonalize two directions that should be independent, such as gender and occupation, while minimizing the change in the embeddings to preserve important semantic meaning from gendered words.
@@ 5.1.6 Discussion and Limitations
Pre-processing mitigations may have limited effectiveness and may rely on questionable assumptions. Data augmentation techniques swap terms using word lists, which can be unscalable and introduce factuality errors (Kumar et al., 2023b). Furthermore, word lists are often limited in length and scope, may depend on proxies (e.g., names as a proxy for gender) that are often tied to other social identities, and utilize word pairs that are not semantically or connotatively equivalent (Devinney et al., 2022). Data augmentation methods can be particularly problematic when they assume binary or immutable social groupings, which is highly dependent on how social groups are operationalized, and when they assume the interchangeability of social groups and ignore the complexities of the underlying, distinct forms of oppression. Data filtering, reweighting, and generation processes may encounter similar challenges, particularly with misrepresentative word lists and proxies for social groups, and may introduce new distribution imbalances into the dataset. That said, instance reweighting opens the door to future work in cost-sensitive learning approaches, and data generation can set a strong standard for careful data curation that can be followed for future datasets.
Instruction tuning also faces a number of challenges. Modified prompting language techniques have been shown to have limited effectiveness. Borchers et al. (2022), for example, find instructions that prompt diversity or gender equality to be unsuccessful for bias removal in outputs. Similarly, Li & Zhang (2023) find similar generated outputs when using biased and unbiased prompts. That said, one benefit of modified prompting language and control tokens is their interpretability, which continuous prompt tuning lacks.
@@ 5.2 In-Training Mitigation
In-training mitigation techniques aim to modify the training procedure to reduce bias. These approaches modify the optimization process by changing the loss function, updating next-word probabilities in training, selectively freezing parameters during fine-tuning, or identifying and removing specific neurons that contribute to harmful outputs. All in-training mitigations change model parameters via gradient-based training updates. We describe each type of in-training mitigation here, and show examples in Figure 8.
Figure 8: Example In-Training Mitigation Techniques (§ 5.2). We illustrate four classes of methods that modify model parameters during training. Architecture modifications change the configuration of the model, such as adding new trainable parameters with adapter modules as done in this example (Lauscher et al., 2021). Loss function modifications introduce a new optimization objective, such as equalizing the embeddings or predicted probabilities of counterfactual tokens or sentences. Selective parameter updates freeze the majority of the weights and only tune a select few during fine-tuning to minimize forgetting of pre-trained language understanding. Filtering model parameters, in contrast, freezes all pre-trained weights and selectively prunes some based on a debiasing objective.
@@ 5.2.1 Architecture Modification
Architecture modifications consider changes to the configuration of a model, including the number, size, and type of layers, encoders, and decoders. For instance, Lauscher et al. (2021) introduce debiasing adapter modules, called ADELE, to mitigate gender bias. The technique is based on modular adapter frameworks (Houlsby et al., 2019) that add new, randomly-initialized layers between the original layers for parameter-efficient fine-tuning; only the injected layers are updated during fine-tuning, while the pre-trained ones remain frozen. This work uses the adapter layers to learn debiasing knowledge by fine-tuning on the BEC-Pro gender bias dataset (Bartl et al., 2020). Ensemble models may also enable bias mitigation. Han et al. (2022a) propose a gated model that takes protected attributes as a secondary input, concatenating the outputs from a shared encoder used by all inputs with the outputs from a demographic-specific encoder, before feeding the combined encodings to the decoder or downstream task.
@@ 5.2.2 Loss Function Modification
Liu et al. (2020)
$\mathcal{R}=\lambda\sum_{(a_{i},a_{j})\in A}\left\|E(a_{i})-E(a_{j})\right\|_{2}$
Yang et al. (2023)
$\mathcal{L}=\Sigma_{i,j\in\{1,\cdots,d\},i<j}JS\left(P^{a_{i}}\|P^{a_{j}}\right)+\lambda KL\left(Q\|P\right)$
Woo et al. (2023)
$\mathcal{R}=\frac{1}{2}\sum_{i\in\{m,f\}}KL\left(E(S_{i})\big{\|}\frac{E(S_{m})+E(S_{f})}{2}\right)-\frac{E(S_{m})^{\top}E(S_{f})}{\|E(S_{m})\|\|E(S_{f})\|}$
Park et al. (2023)
$\mathcal{R}=\sum_{w\in W_{\textrm{stereo}}}\left|\frac{\bm{\mathrm{g}}}{\|\bm{\mathrm{g}}\|}^{\top}w\right|$
Bordia & Bowman (2019)
$\mathcal{R}=\lambda\left\|E(W)B\right\|^{2}_{F}$
Kaneko & Bollegala (2021)
$\mathcal{R}=\sum_{w\in W}\sum_{S\in\mathbb{S}}\sum_{a\in A}\left(\bar{\bm{\mathrm{a}}}_{i}^{\top}E_{i}(w,S)\right)^{2}$
Colombo et al. (2021)
$\mathcal{R}=\lambda I\left(E(X);A\right)$
@@
Gaci et al. (2022)
$\mathcal{L}=\sum_{S\in\mathbb{S}}\sum_{\ell=1}^{L}\sum_{h=1}^{H}\left\|\mathbf{A}_{:\sigma,:\sigma}^{l,h,S,G}-\mathbf{O}_{:\sigma,:\sigma}^{l,h,S,G}\right\|_{2}^{2}$
@@
Attanasio et al. (2022)
$\mathcal{R}=-\lambda\sum_{\ell=1}^{L}\textrm{entropy}(\mathbf{A})^{\ell}$
@@
Qian et al. (2019),
$\mathcal{R}=\lambda\frac{1}{K}\sum_{k=1}^{K}\left|\log\frac{P(a_{i}^{(k)})}{P({a_{j}^{(k)}})}\right|$
@@
Garimella et al. (2021)
$\mathcal{R}(t)=\lambda\left|\log\tfrac{\Sigma_{k=1}^{|A_{i}|}P(A_{i,k})}{\Sigma_{k=1}^{|A_{j}|}P(A_{j,k})}\right|$
Guo et al. (2022)
$\mathcal{L}=\frac{1}{|\mathbb{S}|}\sum_{S\in\mathbb{S}}\sum_{k=1}^{K}JS\left(P({a_{1}}^{(k)}),P({a_{2}}^{(k)}),\cdots,P({a_{m}}^{(k)})\right)$
Garg et al. (2019)
$\mathcal{R}=\lambda\sum_{X\in\mathbb{X}}|z(X_{i})-z(X_{j})|$
He et al. (2022b)
$\mathcal{R}=\lambda\sum_{x\in X}\left\{\begin{matrix}\textrm{\tiny{energy}}_{\textrm{\tiny{task}}}(x)+(\textrm{\tiny{energy}}_{\textrm{\tiny{bias}}}(x)-\tau)&\text{if }\textrm{\tiny{energy}}_{\textrm{\tiny{bias}}}(x)>\tau\\ 0&\text{otherwise}\end{matrix}\right.$
Garimella et al. (2021)
$\mathcal{R}=\sum_{w\in W}\left(e^{\textrm{\tiny{bias}}(w)}\times P(w)\right)$
Embeddings.
@@
In their method GuiDebias, Woo et al. (2023) consider gender stereotype sentences, with a regularization term (Equation 32) to enforce independence between gender groups and the representations of stereotypical masculine $S_{m}$ and feminine $S_{f}$ sentences, given by an encoder $E$. Instead of adding the regularization term to the model’s original loss function, the authors propose an alternative loss to maintain the pre-trained model’s linguistic integrity by preserving non-stereotype sentences.
@@
The second set of work integrates projection-based mitigation techniques (see Section 5.1.5) into the loss function. To mitigate gender stereotypes in occupation terms, Park et al. (2023) introduces a regularization term that orthogonalizes stereotypical word embeddings $w$ and the gender direction $\bm{\mathrm{g}}$ in the embedding space. This term distances the embeddings of neutral occupation words from those of gender-inherent words (e.g., "sister" or "brother"). The gender direction is shown in Equation 33, where $A$ is the set of all gender-inherent feminine-associated $a_{i}$ and masculine-associated $a_{j}$ words, and $E(\cdot)$ computes the embeddings of a model; the regularization term is given by Equation 34, where $W_{\textrm{stereo}}$ is the set of stereotypical embeddings.
$\bm{\mathrm{g}}=\frac{1}{|A|}\sum_{(a_{i},a_{j})\in A}E(a_{j})-E(a_{i})$ (33)
$\mathcal{R}=\sum_{w\in W_{\textrm{stereo}}}\left|\frac{\bm{\mathrm{g}}}{\|\bm{\mathrm{g}}\|}^{\top}w\right|$ (34)
Bordia & Bowman (2019) alternatively obtain the gender subspace $B$ from the singular value decomposition of a stack of vectors representing gender-opposing words (e.g., "man" and "woman"), and minimize the projection of neutral embeddings, denoted $E(W)$, onto that subspace with the regularization term given by Equation 35.
$\mathcal{R}=\lambda\left\|E(W)B\right\|^{2}_{F}$ (35)
Kaneko & Bollegala (2021) similarly encourages hidden representations to be orthogonal to some protected attribute, with a regularization term (Equation 36) summing over the inner products between the embeddings of neutral token $w\in W$ in an input sentence $S\in\mathbb{S}$ and the average embedding $\bar{\bm{\mathrm{a}}}_{i}$ of all encoded sentences containing protected attribute $a\in A$ for an encoder $E$ at layer $i$.
@@
The last set of work considers the mutual information between a social group and the learned representations. Wang et al. (2023) propose a fairness loss over the hidden states of the encoder to minimize the mutual information between the social group of a sentence (e.g., gender) and the sentence semantics (e.g., occupation). Similarly, Colombo et al. (2021) introduce a regularization term (Equation 37) to minimize mutual information $I$ between a random variable $A$ representing a protected attribute and the encoding of an input $X$ with encoder $E$.
@@
Attention.
Some evidence has indicated that the attention layers of a model may be a primary encoder of bias in language models (Jeoung & Diesner, 2022). Gaci et al. (2022) and Attanasio et al. (2022) propose loss functions that modify the distribution of weights in the attention heads of the model to mitigate bias. Gaci et al. (2022) address stereotypes learned in the attention layer of sentence-level encoders by redistributing attention scores, fine-tuning the encoder with an equalization loss that encourages equal attention scores (e.g., to attend to "doctor") with respect to each social group (e.g., "he" and "she"), while minimizing changes to the attention of other words in the sentence. The equalization loss is added as a regularization term to a semantic information preservation term that computes the distance between the original (denoted by $\mathbf{O}$) and fine-tuned models’ attention scores. The equalization loss is given by Equation 38 for a sentence $S\in\mathbb{S}$ and an encoder with $L$ layers, $H$ attention heads, $|\mathbb{G}|$ social groups.
@@
Attanasio et al. (2022) introduce Entropy-based Attention Regularization (EAR). They use the entropy of the attention weights’ distribution to measure the relevance of context words, with a high entropy indicating a wide use of context and a small entropy indicating the reliance on a few select tokens. The authors propose maximizing the entropy of the attention weights to avoid overfitting to stereotypical words in a sentence and to encourage attention to the broader context of the input. Entropy maximization is added as a regularization term to the loss, shown in Equation 39, where $\textrm{entropy}(\mathbf{A})^{\ell}$ is the attention entropy at the $\ell$-th layer.
@@
Predicted token distribution.
Several works propose loss functions that equalize the probability of demographically-associated words in the generated output. Qian et al. (2019), for instance, propose an equalizing objective that encourages demographic words to be predicted with equal probability. They introduce a regularization term comparing the output softmax probabilities $P$ for binary masculine and feminine word pairs, which was adapted by Garimella et al. (2021) for binary race word pairs. The regularization term is shown in Equation 40, for $K$ word pairs consisting of attributes $a_{i}$ and $a_{j}$.
@@
In Auto-Debias, Guo et al. (2022) extend these ideas to non-binary social groups, encouraging the generated output to be independent of social group. The loss, given by Equation 42, calculates the Jensen-Shannon divergence between predicted distributions $P$ conditioned on a prompt $S\in\mathcal{S}$ concatenated with an attribute word $a_{i}$ for $K$ tuples of $m$ attributes (e.g., ("judaism", "christianity", "islam")).
@@
Garg et al. (2019) alternatively consider counterfactual logits, presenting counterfactual logit pairing (CLP). This method encourages the logits of a sentence and its counterfactual (with a protected attribute flipped) to be equal by adding a regularization term to the loss function, given by Equation 43, for the original logit $z(X_{i})$ and its counterfactual $z(X_{j})$.
@@
Zhou et al. (2023) use causal invariance to mitigate gender and racial bias in fine-tuning for a specific downstream task, by treating label-relevant factors to the downstream task as causal, and bias-relevant factors as non-casual. They add a regularization term that enforces equivalent outputs for sentences with the same semantics but different attribute words.
@@
Garimella et al. (2021) assign bias scores bias to all adjectives and adverbs $W$ in the vocabulary to generate a bias penalization regularization term shown in Equation 45.
@@
Dropout.
Instead of proposing a new regularization term, Webster et al. (2020) use dropout regularization during pre-training to reduce stereotypical gendered associations between words. By increasing dropout on the attention weights and hidden activations, the work hypothesizes that the interruption of the attention mechanism disrupts gendered correlations.
Contrastive learning.
Traditional contrastive learning techniques consider the juxtaposition of pairs of unlabeled data to learn similarities or differences within the dataset. As a bias mitigation technique, contrastive loss functions have been adopted to a supervised setting, taking biased-unbiased pairs of sentences and maximizing similarity to the unbiased sentence. The pairs of sentences are often generated by replacing protected attributes with their opposite or an alternative (Cheng et al., 2021; He et al., 2022a; Oh et al., 2022). Cheng et al. (2021)’s FairFil, for instance, trains a network to maximize the mutual information between an original sentence and its counterfactual, while minimizing the mutual information between the outputted embedding and the embeddings of protected attributes. Oh et al. (2022)’s FarconVAE uses a contrastive loss to learn a mapping from the original input to two separate representations in the latent space, one sensitive and one non-sensitive space with respect to some attribute. The non-sensitive representation can be used for downstream predictions. To avoid overfitting to counterfactual pairs, Li et al. (2023) first amplify bias before reducing it with contrastive learning. To amplify bias, they use continuous prompt tuning (by prepending trainable tokens to the start of the input) to increase the difference between sentence pairs. Following other works, the model then trains on a contrastive loss to maximize similarity between the counterfactual sentence pairs.
Other works have proposed alternative contrastive pairs. To debias pre-trained representations, Shen et al. (2022) create positive samples between examples sharing a protected attribute (and, optionally, a class label), and use a negated contrastive loss to discourage the contrasting of instances belonging to different social groups. Khalatbari et al. (2023) propose a contrastive regularization term to reduce toxicity. They learn safe and unsafe distributions from non-toxic and toxic examples, and the contrastive loss pulls the model away from the toxic data distribution while simultaneously pushing it towards the non-toxic data distribution using Jensen-Shannon divergence.

### v3

@@ 5.1.5 Projection-based Mitigation
Ravfogel et al. (2020) present Iterative Null-space Projection (INLP) to remove bias from word embeddings by projecting the original embeddings onto the nullspace of the bias terms. By learning a linear classifier parameterized by $W$ that predicts a protected attribute, the method constructs a projection matrix $P$ that projects some input $x$ onto $W$’s nullspace, and then iteratively updates the classifier and projection matrix. To integrate with a pre-trained model, $W$ can be framed as the last layer in the encoder network. Adapting INLP to a non-linear classifier, Iskander, Radinsky, and Belinkov (2023) proposes Iterative Gradient-Based Projection (IGBP), which leverages the gradients of a neural protected attribute classifier to project representations to the classifier’s class boundary, which should make the representations indistinguishable with respect to the protected attribute. Liang et al. (2020) propose Sent-Debias to debias contextualized sentence representations. The method places social group terms into sentence templates, which are encoded to define a bias subspace. Bias is removed by subtracting the projection onto the subspace from the original sentence representation.
However, removing the concept of gender or any other protected attribute altogether may be too aggressive and eliminate important semantic or grammatical information. To address this, Limisiewicz and Mareček (2022) distinguish a gender bias subspace from the embedding space, without diminishing the semantic information contained in gendered words like pronouns. They use an orthogonal transformation to probe for gender information, and discard latent dimensions corresponding to bias, while keeping dimensions containing grammatical gender information. In their method OSCAR, Dev et al. (2021) also perform less-aggressive bias removal to maintain relevant semantic information. They orthogonalize two directions that should be independent, such as gender and occupation, while minimizing the change in the embeddings to preserve important semantic meaning from gendered words.
@@ 5.1.6 Discussion and Limitations
Pre-processing mitigations may have limited effectiveness and may rely on questionable assumptions. Data augmentation techniques swap terms using word lists, which can be unscalable and introduce factuality errors (Kumar et al., 2023b). Furthermore, word lists are often limited in length and scope, may depend on proxies (e.g., names as a proxy for gender) that are often tied to other social identities, and utilize word pairs that are not semantically or connotatively equivalent (Devinney, Björklund, and Björklund, 2022). Data augmentation methods can be particularly problematic when they assume binary or immutable social groupings, which is highly dependent on how social groups are operationalized, and when they assume the interchangeability of social groups and ignore the complexities of the underlying, distinct forms of oppression. Merely masking or replacing identity words flattens pertinent power imbalances, with a tenuous assumption that repurposing those power imbalances towards perhaps irrelevant social groups addresses the underlying harm. Diminishing the identity of the harmed group is an inadequate patch.
Data filtering, reweighting, and generation processes may encounter similar challenges, particularly with misrepresentative word lists and proxies for social groups, and may introduce new distribution imbalances into the dataset. Data generation derived from crowdsourcing, for instance, may favor majority opinions, as Kim et al. (2022) point out in their creation of an inherently subjective social norm dataset, based on the Social Chemistry dataset that Forbes et al. (2020) acknowledge to represent primarily English-speaking, North American norms.
Instruction tuning also faces a number of challenges. Modified prompting language techniques have been shown to have limited effectiveness. Borchers et al. (2022), for example, find instructions that prompt diversity or gender equality to be unsuccessful for bias removal in outputs. Similarly, Li and Zhang (2023) find similar generated outputs when using biased and unbiased prompts. That said, modified prompting language and control tokens benefits from interpretability, which the continuous prompt tuning lacks.
@@
Despite these limitations, pre-processing techniques also open the door to stronger alternatives. For instance, future work can leverage instance reweighting for cost-sensitive learning approaches when social groups are imbalanced, increasing the weight or error penalty for minority groups. Such approaches can gear downstream training towards macro-averaged optimization that encourages improvement for minority classes. Data generation can set a strong standard for careful data curation that can be followed for future datasets. For example, drawing inspiration from works like Davani, Díaz, and Prabhakaran (2022), Denton et al. (2021), and Fleisig, Abebe, and Klein (2023), future datasets can ensure that the identities, backgrounds, and perspectives of human authors are documented so that the positionality of datasets are not rendered invisible or neutral (Leavy, Siapera, and O’Sullivan, 2021).
@@ 5.2 In-Training Mitigation
In-training mitigation techniques aim to modify the training procedure to reduce bias. These approaches modify the optimization process by changing the loss function, updating next-word probabilities in training, selectively freezing parameters during fine-tuning, or identifying and removing specific neurons that contribute to harmful outputs. All in-training mitigations change model parameters via gradient-based training updates. We describe each type of in-training mitigation here, with examples in Figure 8.
Figure 8: Example In-Training Mitigation Techniques (§ 5.2). We illustrate four classes of methods that modify model parameters during training. Architecture modifications change the configuration of the model, such as adding new trainable parameters with adapter modules as done in this example (Lauscher, Lueken, and Glavaš, 2021). Loss function modifications introduce a new optimization objective, such as equalizing the embeddings or predicted probabilities of counterfactual tokens or sentences. Selective parameter updates freeze the majority of the weights and only tune a select few during fine-tuning to minimize forgetting of pre-trained language understanding. Filtering model parameters, in contrast, freezes all pre-trained weights and selectively prunes some based on a debiasing objective.
@@ 5.2.1 Architecture Modification
Architecture modifications consider changes to the configuration of a model, including the number, size, and type of layers, encoders, and decoders. For instance, Lauscher, Lueken, and Glavaš (2021) introduce debiasing adapter modules, called ADELE, to mitigate gender bias. The technique is based on modular adapter frameworks (Houlsby et al., 2019) that add new, randomly-initialized layers between the original layers for parameter-efficient fine-tuning; only the injected layers are updated during fine-tuning, while the pre-trained ones remain frozen. This work uses the adapter layers to learn debiasing knowledge by fine-tuning on the BEC-Pro gender bias dataset (Bartl, Nissim, and Gatt, 2020). Ensemble models may also enable bias mitigation. Han, Baldwin, and Cohn (2022a) propose a gated model that takes protected attributes as a secondary input, concatenating the outputs from a shared encoder used by all inputs with the outputs from a demographic-specific encoder, before feeding the combined encodings to the decoder or downstream task.
@@ 5.2.2 Loss Function Modification
Liu et al. (2020) $\mathcal{R}=\lambda\sum_{(a_{i},a_{j})\in A}\left\|E(a_{i})-E(a_{j})\right\|_{2}$
Yang et al. (2023) $\mathcal{L}=\Sigma_{i,j\in\{1,\cdots,d\},i<j}JS\left(P^{a_{i}}\|P^{a_{j}}\right)+\lambda KL\left(Q\|P\right)$
Woo et al. (2023) $\mathcal{R}=\frac{1}{2}\sum_{i\in\{m,f\}}KL\left(E(S_{i})\big{\|}\frac{E(S_{m})+E(S_{f})}{2}\right)$
$-\frac{E(S_{m})^{\top}E(S_{f})}{\|E(S_{m})\|\|E(S_{f})\|}$
Park et al. (2023) $\mathcal{R}=\sum_{w\in W_{\textrm{stereo}}}\left|\frac{\bm{\mathrm{v}}_{\textrm{gender}}}{\|\bm{\mathrm{v}}_{\textrm{gender}}\|}^{\top}w\right|$
Bordia and Bowman (2019) $\mathcal{R}=\lambda\left\|E(W)V_{\textrm{gender}}\right\|^{2}_{F}$
Kaneko and Bollegala (2021) $\mathcal{R}=\sum_{w\in W}\sum_{S\in\mathbb{S}}\sum_{a\in A}\left(\bar{\bm{\mathrm{a}}}_{i}^{\top}E_{i}(w,S)\right)^{2}$
Colombo, Piantanida, and Clavel (2021) $\mathcal{R}=\lambda I\left(E(X);A\right)$
@@
Gaci et al. (2022) $\mathcal{L}=\sum_{S\in\mathbb{S}}\sum_{\ell=1}^{L}\sum_{h=1}^{H}\left\|\mathbf{A}_{:\sigma,:\sigma}^{l,h,S,G}-\mathbf{O}_{:\sigma,:\sigma}^{l,h,S,G}\right\|_{2}^{2}$
@@
Attanasio et al. (2022) $\mathcal{R}=-\lambda\sum_{\ell=1}^{L}\textrm{entropy}(\mathbf{A})^{\ell}$
@@
Qian et al. (2019), $\mathcal{R}=\lambda\frac{1}{K}\sum_{k=1}^{K}\left|\log\frac{P(a_{i}^{(k)})}{P({a_{j}^{(k)}})}\right|$
@@
Garimella et al. (2021) $\mathcal{R}(t)=\lambda\left|\log\tfrac{\Sigma_{k=1}^{|A_{i}|}P(A_{i,k})}{\Sigma_{k=1}^{|A_{j}|}P(A_{j,k})}\right|$
Guo, Yang, and Abbasi (2022) $\mathcal{L}=\frac{1}{|\mathbb{S}|}\sum_{S\in\mathbb{S}}\sum_{k=1}^{K}JS\left(P({a_{1}}^{(k)}),P({a_{2}}^{(k)}),\cdots,P({a_{m}}^{(k)})\right)$
Garg et al. (2019) $\mathcal{R}=\lambda\sum_{X\in\mathbb{X}}|z(X_{i})-z(X_{j})|$
He et al. (2022b) $\mathcal{R}=\lambda\sum_{x\in X}\left\{\begin{matrix}\textrm{\tiny{energy}}_{\textrm{\tiny{task}}}(x)+(\textrm{\tiny{energy}}_{\textrm{\tiny{bias}}}(x)-\tau)&\text{if }\textrm{\tiny{energy}}_{\textrm{\tiny{bias}}}(x)>\tau\\ 0&\text{otherwise}\end{matrix}\right.$
Garimella et al. (2021) $\mathcal{R}=\sum_{w\in W}\left(e^{\textrm{\tiny{bias}}(w)}\times P(w)\right)$
Embeddings
@@
In their method GuiDebias, Woo et al. (2023) consider gender stereotype sentences, with a regularization term (Equation 32) to enforce independence between gender groups and the representations of stereotypical masculine $S_{m}$ and feminine $S_{f}$ sentences, given by the hidden representations $E$ in the last layer. Instead of adding the regularization term to the model’s original loss function, the authors propose an alternative loss to maintain the pre-trained model’s linguistic integrity by preserving non-stereotype sentences.
@@
The second set of work integrates projection-based mitigation techniques (see Section 5.1.5) into the loss function. To mitigate gender stereotypes in occupation terms, Park et al. (2023) introduces a regularization term that orthogonalizes stereotypical word embeddings $w$ and the gender direction $\bm{\mathrm{v}}_{\textrm{gender}}$ in the embedding space. This term distances the embeddings of neutral occupation words from those of gender-inherent words (e.g., "sister" or "brother"). The gender direction is shown in Equation 33, where $A$ is the set of all gender-inherent feminine-associated $a_{i}$ and masculine-associated $a_{j}$ words, and $E(\cdot)$ computes the embeddings of a model; the regularization term is given by Equation 34, where $W_{\textrm{stereo}}$ is the set of stereotypical embeddings.
$\bm{\mathrm{v}}_{\textrm{gender}}=\frac{1}{|A|}\sum_{(a_{i},a_{j})\in A}E(a_{j})-E(a_{i})$ (33)
$\mathcal{R}=\sum_{w\in W_{\textrm{stereo}}}\left|\frac{\bm{\mathrm{v}}_{\textrm{gender}}}{\|\bm{\mathrm{v}}_{\textrm{gender}}\|}^{\top}w\right|$ (34)
Bordia and Bowman (2019) alternatively obtain the gender subspace $B$ from the singular value decomposition of a stack of vectors representing gender-opposing words (e.g., "man" and "woman"), and minimize the squared Frobenius norm of the projection of neutral embeddings, denoted $E(W)$, onto that subspace with the regularization term given by Equation 35.
$\mathcal{R}=\lambda\left\|E(W)V_{\textrm{gender}}\right\|^{2}_{F}$ (35)
Kaneko and Bollegala (2021) similarly encourages hidden representations to be orthogonal to some protected attribute, with a regularization term (Equation 36) summing over the inner products between the embeddings of neutral token $w\in W$ in an input sentence $S\in\mathbb{S}$ and the average embedding $\bar{\bm{\mathrm{a}}}_{i}$ of all encoded sentences containing protected attribute $a\in A$ for an embedding $E$ at layer $i$.
@@
The last set of work considers the mutual information between a social group and the learned representations. Wang, Cheng, and Henao (2023) propose a fairness loss over the hidden states of the encoder to minimize the mutual information between the social group of a sentence (e.g., gender) and the sentence semantics (e.g., occupation). Similarly, Colombo, Piantanida, and Clavel (2021) introduce a regularization term (Equation 37) to minimize mutual information $I$ between a random variable $A$ representing a protected attribute and the encoding of an input $X$ with hidden representation $E$.
@@
Attention
Some evidence has indicated that the attention layers of a model may be a primary encoder of bias in language models (Jeoung and Diesner, 2022). Gaci et al. (2022) and Attanasio et al. (2022) propose loss functions that modify the distribution of weights in the attention heads of the model to mitigate bias. Gaci et al. (2022) address stereotypes learned in the attention layer of sentence-level encoders by redistributing attention scores, fine-tuning the encoder with an equalization loss that encourages equal attention scores (e.g., to attend to "doctor") with respect to each social group (e.g., "he" and "she"), while minimizing changes to the attention of other words in the sentence. The equalization loss is added as a regularization term to a semantic information preservation term that computes the distance between the original (denoted by $\mathbf{O}$) and fine-tuned models’ attention scores. The equalization loss is given by Equation 38 for a sentence $S\in\mathbb{S}$ and an encoder with $L$ layers, $H$ attention heads, $|\mathbb{G}|$ social groups.
@@
Attanasio et al. (2022) introduce Entropy-based Attention Regularization (EAR), following Ousidhoum et al. (2021)’s observation that models may overfit to identity words and thus overrely on identity terms in a sentence in prediction tasks. They use the entropy of the attention weights’ distribution to measure the relevance of context words, with a high entropy indicating a wide use of context and a small entropy indicating the reliance on a few select tokens. The authors propose maximizing the entropy of the attention weights to encourage attention to the broader context of the input. Entropy maximization is added as a regularization term to the loss, shown in Equation 39, where $\textrm{entropy}(\mathbf{A})^{\ell}$ is the attention entropy at the $\ell$-th layer.
@@
Predicted token distribution
Several works propose loss functions that equalize the probability of demographically-associated words in the generated output. Qian et al. (2019), for instance, propose an equalizing objective that encourages demographic words to be predicted with equal probability. They introduce a regularization term comparing the output softmax probabilities $P$ for binary masculine and feminine words pairs, which was adapted by Garimella et al. (2021) for binary race word pairs. The regularization term is shown in Equation 40, for $K$ word pairs consisting of attributes $a_{i}$ and $a_{j}$.
@@
In Auto-Debias, Guo, Yang, and Abbasi (2022) extend these ideas to non-binary social groups, encouraging the generated output to be independent of social group. The loss, given by Equation 42, calculates the Jensen-Shannon divergence between predicted distributions $P$ conditioned on a prompt $S\in\mathcal{S}$ concatenated with an attribute word $a_{i}$ for $K$ tuples of $m$ attributes (e.g., ("judaism," "christianity," "islam")).
@@
Garg et al. (2019) alternatively consider counterfactual logits, presenting counterfactual logit pairing (CLP). This method encourages the logits of a sentence and its counterfactual to be equal by adding a regularization term to the loss function, given by Equation 43, for the original logit $z(X_{i})$ and its counterfactual $z(X_{j})$.
@@
Zhou et al. (2023) use causal invariance to mitigate gender and racial bias in fine-tuning, by treating label-relevant factors to the downstream task as causal, and bias-relevant factors as non-casual. They add a regularization term to enforce equivalent outputs for sentences with the same semantics but different attribute words.
@@
Garimella et al. (2021) assign bias scores to all adjectives and adverbs $W$ in the vocabulary to generate a bias penalization regularization term shown in Equation 45.
@@
Dropout
Instead of proposing a new regularization term, Webster et al. (2020) use dropout (Srivastava et al., 2014) during pre-training to reduce stereotypical gendered associations between words. By increasing dropout on the attention weights and hidden activations, the work hypothesizes that the interruption of the attention mechanism disrupts gendered correlations.
Contrastive learning
Traditional contrastive learning techniques consider the juxtaposition of pairs of unlabeled data to learn similarity or differences within the dataset. As a bias mitigation technique, contrastive loss functions have been adopted to a supervised setting, taking biased-unbiased pairs of sentences and maximizing similarity to the unbiased sentence. The pairs of sentences are often generated by replacing protected attributes with their opposite or an alternative (Cheng et al., 2021; He et al., 2022a; Oh et al., 2022). Cheng et al. (2021)’s FairFil, for instance, trains a network to maximize the mutual information between an original sentence and its counterfactual, while minimizing the mutual information between the outputted embedding and the embeddings of protected attributes. Oh et al. (2022)’s FarconVAE uses a contrastive loss to learn a mapping from the original input to two separate representations in the latent space, one sensitive and one non-sensitive space with respect to some attribute such as gender. The non-sensitive representation can be used for downstream predictions. To avoid overfitting to counterfactual pairs, Li et al. (2023) first amplify bias before reducing it with contrastive learning. To amplify bias, they use continuous prompt tuning (by prepending trainable tokens to the start of the input) to increase the difference between sentence pairs. The model then trains on a contrastive loss to maximize similarity between the counterfactual sentence pairs.
Other works have proposed alternative contrastive pairs. To debias pre-trained representations, Shen et al. (2022) create positive samples between examples sharing a protected attribute (and, optionally, a class label), and use a negated contrastive loss to discourage the contrasting of instances belonging to different social groups. Khalatbari et al. (2023) propose a contrastive regularization term to reduce toxicity. They learn distributions from non-toxic and toxic examples, and the contrastive loss pulls the model away from the toxic data distribution while simultaneously pushing it towards the non-toxic data distribution using Jensen-Shannon divergence.